### v1

# Privacy of the last iterate in cyclically-sampled DP-SGD on nonconvex composite losses

Abstract
Differentially private stochastic gradient descent (DP-SGD) refers to a family of optimization algorithms that provide a guaranteed level of differential privacy (DP) through DP accounting techniques. However, current accounting techniques make assumptions that diverge significantly from practical DP-SGD implementations. For example, they may assume the loss function is Lipschitz continuous and convex, sample the batches randomly with replacement, or omit the gradient clipping step.
In this work, we analyze the most commonly used variant of DP-SGD, in which we sample batches cyclically with replacement, perform gradient clipping, and only release the last DP-SGD iterate. More specifically — without assuming convexity, smoothness, or Lipschitz continuity of the loss function — we establish new Rényi differential privacy (RDP) bounds for the last DP-SGD iterate under the mild assumption that (i) the DP-SGD stepsize is small relative to the topological constants in the loss function, and (ii) the loss function is weakly-convex. Moreover, we show that our bounds converge to previously established convex bounds when the weak-convexity parameter of the objective function approaches zero. In the case of non-Lipschitz smooth loss functions, we provide a weaker bound that scales well in terms of the number of DP-SGD iterations.
1 Introduction
Tight privacy guarantees for DP-SGD are challenging to achieve in real-world implementations because existing approaches rely on simplifying assumptions. Some common ones include convexity and Lipschitz continuity of the loss function, randomly sampling the batches, and omitting the ubiquitous gradient clipping step. More crucially, existing practical methods assume that the weights of a trained model are always released at every DP-SGD iteration.
In this work, we develop a family of RDP bounds on the last iterate of DP-SGD. Moreover, these bounds are novel in that they:
  • (i)
    are developed under a DP-SGD gradient clipping step without assuming that the objective function is Lipschitz continuous;
  • (ii)
    require substantially weaker assumptions compared to existing works;
  • (iii)
    are parameterized by a weak convexity parameter $m\geq 0$, for which the bounds smoothly converge to similar ones in the convex case as $m\to 0$ when the loss function has a Lipschitz-smooth component; and
  • (iv)
    can be made smaller by decreasing the SGD stepsize, increasing the standard deviation in the DP-SGD Gaussian noise, or increasing the batch size.
Problem definition. More precisely, we develop RDP bounds for the last iterates of DP-SGD applied to the composite optimization problem
$\min_{x\in\mathbb{R}^{n}}\left\{\phi(x):=\frac{1}{k}\sum_{i=1}^{k}f_{i}(x)+h(x)\right\}$ (1)
where $h$ is convex and proper lower-semicontinuous and $f_{i}$ is continuously differentiable on the domain of $h$. Notice that the assumption on $h$ encapsulates (i) common nonsmooth regularization functions such as the $\ell_{1}$-norm $\|\cdot\|_{1}$, nuclear matrix norm $\|\cdot\|_{*}$, and elastic net regularizer and (ii) indicator functions on possibly unbounded closed convex sets. A common setting in the wild is when $(1/k)\sum_{i=1}^{k}f_{i}(x)$ corresponds to a softmax cross-entropy loss function and $h(x)$ corresponds to an $\ell_{1}$- or $\ell_{2}$-regularization function.
Types of bounds. Each of the established bounds is of the form
$D_{\alpha}(X_{T}\|X_{T}^{\prime})\preceq\frac{\alpha\cdot{\cal{B}}_{\lambda}(C,b,T,\ell)}{\sigma^{2}}$ (2)
for some finite function $\cal B_{\lambda}(\cdot)$ and $\lambda>0$, where (i) $D_{\alpha}(X\|X^{\prime})$ denotes the $\alpha$-Rényi divergence between random variables $X$ and $X^{\prime}$, (ii) $X_{T}$ and $X_{T}^{\prime}$ are the last DP-SGD iterates on neighboring datasets, (iii) and the paramaters $\sigma$, $\lambda$, $C$, $b$, $T$, and $\ell$ are DP-SGD’s standard deviation, stepsize, $\ell_{2}$ clip norm, batch size, number of steps, and number of steps per dataset pass, respectively.
More specifically, we target three regimes. First, without any additional assumptions, we obtain (2) with $B_{\lambda}(C,b,T,\ell)=T(\lambda C)^{2}$. Second, assuming that the DP-SGD iterates are contained within an $\ell_{2}$ ball of diameter $d_{h}$, we obtain (2) with $B_{\lambda}(C,b,T,\ell)=(d_{h}+\lambda C/b)^{2}$ for small enough $\lambda$. Third, assuming that each $\nabla f_{i}$ is Lipschitz continuous, we obtain (2) with
$B_{\lambda}(C,b,T,\ell)=\frac{T}{\ell}\left(\frac{L^{2\ell}_{\lambda}}{\sum_{i=1}^{\ell}L^{2i}_{\lambda}}\right)\left(\frac{\lambda C}{b}\right)^{2},$
for some finite $L_{\lambda}\leq\sqrt{3}$ and small enough $\lambda$. Furthermore, we show that the bound obtained in the third regime converges to previously established bounds in the convex setting when each $f_{i}$ becomes increasingly more convex (specifically when $m\to 0$ in contribution (iii)).
Lipschitz continuity. It is worth re-emphasizing that we do not require each $f_{i}$ in (1) to be Lipschitz continuous in order to bound $\nabla f_{i}$ (see, for example, [8, 2, 7] which do require this assumption). As a consequence, our analysis is applicable to a substantially wider class of non-Lipschitzian objective functions, including (a) global or piecewise quadratic functions on $\mathbb{R}^{n}$, (b) the function $f(x)=x\log x$ on $1,\infty)$ and its multivariate extensions, and (c) double integrals of bounded sigmoidal functions, e.g., $f(x)=\int\log\cosh x\,dx=\int\int\tanh x\,dx$, and their multivariate extensions.
1.1 Mathematical techniques
Weak convexity. We generalize existing analyses in the convex and twice-differentiable setting (see, for example, [8, 7]) to the weakly-convex (and possibly nonsmooth) composite setting. Crucially, our approach avoids assuming the existence of the Hessian for the smooth part of the objective function (see, for example, Lemma 3.2 and Proposition 3.4). This generalization is achieved by leveraging recently developed optimal transport techniques [8] and a curvature-based characterization of nonconvex functions described in Section 3. This characterization specifically enables us to show how our bounds converge to similar ones in the convex setting by appropriately scaling a weak convexity parameter.
Upper curvature. To facilitate a more general analysis, we assume that there exists a constant $M>0$ such that each $f_{i}+M\|\cdot\|^{2}/2$ is convex (see Section 3 for an in-depth discussion). This assumption, together with weak-convexity, is mild in that it is implied by assuming each $\nabla f_{i}$ is $L$-Lipschitz continuous for a large enough $L$. On the other hand, it is essential for establishing that DP-SGD is a Lipschitz operator for small enough stepsize (see Proposition 3.4) and is classical in the analysis of first-order optimization algorithms.
Proximal operator. To analyze the effect of gradient clipping, we utilize important properties about proximal operators, such as strict non-expansiveness and uniqueness for closed convex functions. These properties are specifically applied when we view both the projection and clipping steps of DP-SGD as instances of a particular proximal evaluation. Conversely, our analysis may be applied to more general forms of DP-SGD where the gradient clipping step is replaced by any sequence of proximal evaluations (see, for example, Proposition 3.4).
Shifted Rényi divergence. We generalize the approach of shifted Rényi divergences in [8, 2] from nonexpansive [^3] In [8, 2], the authors refer to the relevant operators as contractions even though they are only $1$-Lipschitz, i.e., nonexpansive. operators to $L$-Lipschitz operators for $L>1$. More precisely, while we still sum a sequence of parameterized shifted Rényi divergence bounds, we introduce more parameters compared to previous works to deal with the issue of nonconvexity. Furthermore, we show that in the setting of one dataset pass, the optimal choice of parameters is obtained by solving a related quadratic programming problem (see Appendix B). For the case of multiple dataset passes, we use the optimal parameter in the single dataset setting together with some nice properties of cyclic DP-SGD batch updates (see Proposition 2.2).
1.2 Related works
Recent approaches [8, 7, 6, 2] directly bound the Rényi divergence of the last iterate, but these approaches make strong assumptions that limit their practical viability. In particular, papers [8, 7] only consider the convex setting. Furthermore, they assume Lispchitz continuity of the loss function, which consequently bounds the loss gradients. While paper [6] studies multiple epochs or passes over the data, their results only apply to the smooth and strongly convex setting. Similarly, results in paper[2] consider only convex Lipschitz smooth functions.
Paper [3] analyzes the privacy loss dynamics for nonconvex functions, but their analysis differs from ours in two ways. First, they assume that their DP-SGD batches are obtained by Poisson sampling or sampling without replacement. Second, their results require numerically solving a minimization problem that can be hard for certain parameters and thus their bound is unclear to converge. Tangentially, prior works on DP-SGD accounting often rely on loose bounds that assume the release of intermediate updates [4, 1]. These works rely in differential privacy advanced composition results [9], resulting in noise with standard deviation that scales as $\sqrt{T}$ [4, 1].
In Table 1.2, we compare our bounds against those established using the “Privacy Amplification by Iteration” (PABI) argument in [8] for multi-epoch noisy-SGD under similar smoothness and scaling assumptions. Note that the multi-epoch noisy-SGD algorithm in [8, Algorithm] only considers the case of $E=\ell$ and does not consider batched gradients. As a consequence of the latter, its corresponding RDP upper bound does not depend on the batch size $b$.
Convex Nonconvex $(d_{h}=\infty)$ Nonconvex $(d_{h}<\infty)$ Assumptions
PABI[8] $O\left(\dfrac{\alpha E}{\ell}\left[\dfrac{\lambda Q}{\sigma}\right]^{2}\right)$ - - $f$ is $Q$-Lipschitz; no clipping; no batching
This work $O\left(\dfrac{\alpha E}{\ell}\left[\dfrac{\lambda C}{b\sigma}\right]^{2}\right)$ $O\left(\dfrac{\alpha EL_{\lambda}^{2\ell}}{\sum_{i=0}^{\ell-1}L_{\lambda}^{2i}}\left[\dfrac{\lambda C}{b\sigma}\right]^{2}\right)$ $O\left(\dfrac{\alpha}{\sigma^{2}}\left[L_{\lambda}d_{h}+\dfrac{\lambda C}{b}\right]^{2}\right)$ clipping with an $\ell_{2}$ bound of $C$; batch size $b$
Table 1: Asymptotic $\alpha$-Rényi differential privacy upper bounds of the last iterate after $T$ iterations of DP-SGD, when $f_{i}$ is Lipschitz smooth and weakly-convex for $i\in[k]$, $\lambda$ is the stepsize, $L_{\lambda}=\Theta(\sqrt{1+\lambda m})$ for a weak-convexity parameter $m$ (see (9)), $\sigma$ is the standard deviation of the Gaussian noise, $\ell$ is the number of iterations in one dataset pass, $E$ is the number of dataset passes, and $d_{h}$ is the diameter of the domain of $h$.
1.3 Organization
Section 2 presents some important properties about a specific family of randomized operators. Section 3 states some additional assumptions on (1) and presents some crucial properties for an SGD-like update. Section 4 presents the main RDP bounds of this work. Finally, the appendices at the end of this paper contain technical proofs and extended discussions about the main results.
1.4 Notation
Let $n]=\{1,\ldots,n\}$ for any positive integer $n$. Let $\mathbb{R}$ denote the set of real numbers and $\mathbb{R}^{n}=\mathbb{R}\times\cdots\times\mathbb{R}$ denote the $n$-fold Cartesian product of $\mathbb{R}$. Let $(\langle\cdot,\cdot\rangle,\mathbb{R}^{n})$ denote a Euclidean space over $\mathbb{R}^{n}$ and denote $\|\cdot\|:=\sqrt{\langle\cdot,\cdot\rangle}$ to be the induced norm. The domain of a function $\phi:\mathbb{R}^{n}\mapsto(-\infty,\infty]$ is $\operatorname*{dom}\phi:=\{z\in\mathbb{R}^{n}:\phi(z)<\infty\}$. The function $\phi$ is said to be proper if $\operatorname*{dom}\phi\neq\emptyset$. A function $\phi:\mathbb{R}^{n}\mapsto(-\infty,\infty]$ is said to be lower semicontinuous if $\liminf_{x\to x_{0}}\phi(x)\geq\phi(x_{0})$. The set of proper, lower semicontinuous, convex functions over $\mathbb{R}^{n}$ is denoted by $\overline{{\rm Conv}}\,(\mathbb{R}^{n})$. An operator ${\cal T}:\mathbb{R}^{n}\mapsto\mathbb{R}^{m}$ is said to be $L$-Lipschitz if $\|{\cal T}(x)-{\cal T}(y)\|\leq L\|x-y\|$ for all $x,y\in\mathbb{R}^{n}$, and $\cal T$} is said to be nonexpansive if it is 1-Lipschitz. The proximal operator ${\rm prox}_{\psi}:\mathbb{R}^{n}\mapsto\mathbb{R}^{n}$ for $\psi\in\overline{{\rm Conv}}\,(\mathbb{R}^{n})$ is defined as
$\operatorname{prox}_{\psi}(z_{0})=\operatorname*{argmin}_{z\in\mathbb{R}^{n}}\left\{\psi(z)+\frac{1}{2}\|z-z_{0}\|^{2}\right\}\quad\forall z_{0}\in\mathbb{R}^{n},$
and it is well-known that ${\rm prox}_{\psi}(\cdot)$ is nonexpansive (see, for example [5, Theorem 6.42]).
The $\infty$-Wasserstein metric ${\cal W}_{\infty}(\mu,\nu)$ is the smallest real number $w$ such that for $X\sim\mu$ and $Y\sim\nu$, there is a joint distribution on $(X,Y)$ where $\|X-Y\|\leq w$ almost surely, i.e., ${\cal W}_{\infty}(\mu,\nu)=\inf_{\gamma\sim\Gamma(\mu,\nu)}\operatorname*{ess\ sup}_{(x,y)\sim\gamma}\|x-y\|$, where $\Gamma(\mu,\nu)$ is the collection of measures on $\mathbb{R}^{n}\times\mathbb{R}^{n}$ with first and second marginals $\mu$ and $\nu$, respectively. For any probability distributions $\mu$ and $\nu$ with $\nu\ll\mu$, the Rényi divergence of order $\alpha\in(1,\infty)$ is
$D_{\alpha}(\mu\|\nu)=\frac{1}{\alpha-1}\log\int\left[\frac{\mu(x)}{\nu(x)}\right]^{\alpha}\nu(x)\,dx$
where we take the convention that $0/0=0$. For $\nu\not\ll\mu$, we define $D_{\alpha}(\mu\|\nu)=\infty$. For parameters $\tau\geq 0$ and $\alpha\geq 1$, the shifted Rényi divergence is
$D_{\alpha}^{(\tau)}(\mu\|\nu):=\inf_{\gamma}\left\{D_{\alpha}(\gamma\|\nu):{\cal W}_{\infty}(\mu,\gamma)\leq\tau\right\}$ (3)
for any probability distributions $\mu$ and $\nu$ over $\mathbb{R}^{n}$. Given random variables $X\sim\mu$ and $Y\sim\nu$, we denote $D_{\alpha}(X\|Y)=D_{\alpha}(\mu\|\nu)$ and $D_{\alpha}^{(\tau)}(X\|Y)=D_{\alpha}^{(\tau)}(\mu\|\nu)$.
We consider the swap model for differential privacy. We say two datasets $S$ and $S^{\prime}$ are neighbors, denoted as $S\sim S^{\prime}$, if $S^{\prime}$ can be obtained by swapping one record. A randomized algorithm $\cal A$ is said to be $(\alpha,\varepsilon)$-Rényi differentially private (RDP) if, for every pair of neighboring datasets $S$ and $S^{\prime}$ in the domain of $\cal A$, we have $D_{\alpha}({\cal A}(S)\|{\cal A}(S^{\prime}))\leq\varepsilon$.
2 Randomized proximal Lipschitz operators
This section gives some crucial properties about randomized proximal Lipschitz operators. More specifically, it establishes several RDP bounds based on the closeness of neighboring operators.
We first bound the shifted Rényi divergence of a randomized proximal Lipschitz operator. The proof of this result is a simple extension of the argument in [8, Theorem 22] from $1$-Lipschitz operators to $L$-Lipschitz operators with additive residuals. Nevertheless, we give the full argument in Appendix A.1 for completeness.
Lemma 2.1.
For some $L,\zeta\geq 0$, suppose $\phi^{\prime}$ and $\phi$ satisfy
$\sup_{u}\|\phi^{\prime}(u)-\phi(u)\|\leq s,\quad\|\Phi(x)-\Phi(y)\|\leq L\|x-y\|+\zeta,\quad\forall\Phi\in\{\phi,\phi^{\prime}\},$ (4)
for any $x,y\in\operatorname*{dom}\phi\cap\operatorname*{dom}\phi^{\prime}$. Moreover, let $Z\sim{\cal N}(0,\sigma^{2}I)$ and $\psi\in\overline{{\rm Conv}}\,(\mathbb{R}^{n})$. Then, for any scalars $a,\tau\geq 0$ and $\alpha\in(1,\infty)$ satisfying $L\tau+\zeta+s-a\geq 0,$ and random variables $Y$ and $Y^{\prime}$, it holds that
$D_{\alpha}^{(L\tau+\zeta+s-a)}(\operatorname{prox}_{\psi}(\phi(Y)+Z)\|\operatorname{prox}_{\psi}(\phi^{\prime}(Y^{\prime})+Z))\leq D_{\alpha}^{(\tau)}(Y\|Y^{\prime})+\frac{\alpha a^{2}}{2\sigma^{2}}.$ (5)
Note that the second inequality in (4) is equivalent to $\Phi$ being $L$-Lipschitz continuous when $\zeta=0$, and the conditions in (4) need to only hold on $\operatorname*{dom}\phi\cap\operatorname*{dom}\phi^{\prime}$.
We now apply (5) to a sequence of points generated by the randomized proximal update $Y\xleftarrow{}\operatorname{prox}_{\psi}(\phi(Y)+Z)$, under different assumptions on $\zeta$ and $\tau$. This particular result generalizes the one in [8], which only considers the case of $L=1$ and $\zeta=0$. For conciseness, we leave the full proof in Appendix A.2.
Before proceeding, we define the following useful quantities:
$\theta_{L}(0)=0,\quad\theta_{L}(s):=\frac{L^{2(s-1)}}{\sum_{j=0}^{s-1}L^{2j}}\quad\forall s\geq 1.$ (6)
Proposition 2.2.
Let $L,\zeta\geq 0$, $T\geq 1$, and $\ell\in[T]$ be fixed. Given $\psi\in\overline{{\rm Conv}}\,(\mathbb{R}^{n})$, suppose $\{\phi_{t}\}$, $\{\phi_{t}^{\prime}\}$, and $\bar{s}>0$ satisfy (4) with
$\phi=\phi_{t},\quad\phi^{\prime}=\phi_{t}^{\prime},\quad s=\begin{cases}\bar{s},&t=1\bmod\ell,\\ 0,&\text{otherwise},\end{cases}\quad\forall t\in[T].$
Moreover, given $Y_{0},Y_{0}^{\prime}\in\operatorname*{dom}\psi$, let $Z_{t}\sim{\cal N}(0,\sigma^{2}I)$, and define the random variables
$\begin{gathered}Y_{t}:=\operatorname{prox}_{\psi}(\phi_{t}(Y_{t-1})+Z_{t}),\quad Y_{t}^{\prime}:=\operatorname{prox}_{\psi}(\phi_{t}^{\prime}(Y_{t-1}^{\prime})+Z_{t}),\quad\forall t\geq 1\end{gathered}$
Then, for any $\tau\geq 0$ and $E=\lfloor T/\ell\rfloor$, the following statements hold:
  • (a)
    if $\zeta=0$, then
    $D_{\alpha}(Y_{T}\|Y_{T}^{\prime})-D_{\alpha}^{(\tau)}(Y_{0}\|Y_{0}^{\prime})$
    $\leq\frac{\alpha}{2\sigma^{2}}\left[(L\tau+\bar{s})^{2}\theta_{L}(\ell)+\bar{s}^{2}\left\{(E-1)\theta_{L}(\ell)+\theta_{L}(T-E\ell)\right\}\right].$ (7)
  • (b)
    if $\tau=0$ and $\zeta=\bar{s}$, then
    $D_{\alpha}(Y_{T}\|Y_{T}^{\prime})-D_{\alpha}(Y_{0}\|Y_{0}^{\prime})\leq 2\alpha T\left(\frac{\zeta}{\sigma}\right)^{2}.$ (8)
Some remarks about Proposition 2.2 are in order. First, part (a) shows that if $\phi_{t}$ and $\phi_{t}^{\prime}$ only differ at $t=1$, then $D_{\alpha}(Y_{T}\|Y_{T}^{\prime})$ is finite for any $T$. Second, part (a) also shows that if $\phi_{t}$ and $\phi_{t}^{\prime}$ differ cyclically for a cycle length of $\ell$, then the divergence between $Y_{T}$ and $Y_{T}^{\prime}$ grows linearly with the number of cycles $E$. Third, part (b) gives a bound that is independent of $L$. Finally, both of the bounds in parts (a) and (b) can be viewed as Rényi divergences between the Gaussian random variables ${\cal N}(0,\sigma^{2}I)$ and ${\cal N}(\mu,\sigma^{2}I)$ for different values of $\mu$.
In Appendix B, we give a detailed discussion of how the residuals $a$ from Lemma 2.1 are chosen to prove Proposition 2.2(a). In particular, we prove that the chosen residuals yield the tightest RDP bound that can achieved by repeatedly applying (5).
3 Curvature assumptions and operator analysis
This section states some curvature assumptions for future reference and develops the properties of an SGD-like update.
We start by adding assumptions on (1) for use in later subsections. Given $h:\mathbb{R}^{n}\mapsto(-\infty,\infty]$ and $f_{i}:\operatorname*{dom}h\mapsto\mathbb{R}$ for $i\in[k]$, define the assumptions:
  • (A1)
    $h\in\overline{{\rm Conv}}\,(\mathbb{R}^{n})$;
  • (A2)
    there exists $m,M\geq 0$ such that for $i\in[k]$ the function $f_{i}$ is differentiable on an open set containing $\operatorname*{dom}h$ and
    $\frac{m}{2}\|x-y\|^{2}\leq f_{i}(x)-f_{i}(y)-\left\langle\nabla f_{i}(y),x-y\right\rangle\leq\frac{M}{2}\|x-y\|^{2}\quad\forall x,y\in\operatorname*{dom}h.$ (9)
We now give a few remarks about (A1)–(A2). First, (A1) is necessary to ensure that $\operatorname{prox}_{h}(\cdot)$ is well-defined. Second, it can be shown [^4] See, for example, [5, 11] and [10, Proposition 2.1.55]. that (A2) is equivalent to the assumption that $\nabla f$ is $\max\{m,M\}$-Lipschitz continuous when $m=M$. Third, the lower bound in (9) is equivalent to assuming that $f_{i}(\cdot)+m\|\cdot\|^{2}/2$ is convex and, hence, if $m=0$ then $f_{i}$ is convex. The parameter $m$ is often called a weak-convexity parameter of $f_{i}$. Fourth, using symmetry arguments and the third remark, if $M=0$ then $f_{i}$ is concave. Finally, the third remark motivates why we choose two parameters, $m$ and $M$, in (9). Specifically, we use $m$ (resp. $M$) to develop results that can be described in terms of the level of convexity (resp. concavity) of the problem.
We now develop the some properties of an SGD-like update. Given $\{q_{i}\}\subseteq\overline{{\rm Conv}}\,(\mathbb{R}^{n})$ with $\operatorname*{dom}q_{i}\subseteq\operatorname*{dom}h$ and $B\subseteq[k]$, define the prox-linear operator
${\cal A}_{\lambda}(x)={\cal A}_{\lambda}(x,\{f_{i}\},\{q_{i}\}):=x-\frac{\lambda}{|B|}\sum_{i\in B}\operatorname{prox}_{q_{i}}(\nabla f_{i}(x)).$ (10)
Clearly, when $\operatorname{prox}_{q_{i}}(y)=y$ for every $y\in\mathbb{R}^{n}$, the above update corresponds to a SGD step applied to $\min_{x}\sum_{i=1}^{k}f_{i}(x)$ with stepsize $\lambda$ and starting point $x$. Moreover, while it is straightforward to show that ${\cal A}_{\lambda}(\cdot)$ is $(1+\lambda\max\{m,M\})$-Lipschitz continuous when $\{f_{i}\}$ satisfies (A2) [^5] This follows by using the triangle inequality and the non-expansiveness of $\operatorname{prox}_{q_{i}}(\cdot)$., we prove that the Lipschitz constant can be improved to $\Theta(\sqrt{1+\lambda m})$ when $\lambda$ is small. Notice, in particular, that the former bound does not converge to one when $m\to 0$, e.g., when $f_{i}$ becomes more convex, while the latter does.
To start, we recall the following well-known bound from convex analysis. Its proof can be found, for example, in [5, Theorem 5.8(iv)].
Lemma 3.1.
Let $F:\operatorname*{dom}h\mapsto\mathbb{R}$ be convex and differentiable. Then $F$ satisfies
$F(x)-F(y)-\left\langle\nabla F(y),x-y\right\rangle\leq\frac{L}{2}\|x-y\|^{2}\quad\forall x,y\in\operatorname*{dom}h$ (11)
if and only if
$\left\langle\nabla F(x)-\nabla F(y),x-y\right\rangle\geq\frac{1}{L}\|\nabla F(x)-\nabla F(y)\|^{2}\quad\forall x,y\in\operatorname*{dom}h.$
We next give a technical bound on $f_{i}$, which generalizes the co-coercivity of convex functions to weakly-convex functions.
Lemma 3.2.
For any $x,y\in\operatorname*{dom}h$ and $f_{i}$ satisfying (A2), it holds that
$\left\langle\nabla f_{i}(x)-\nabla f_{i}(y),x-y\right\rangle\geq-\left[m+\frac{1}{2(M+m)}\right]\|x-y\|^{2}+\frac{1}{2(M+m)}\|\nabla f_{i}(x)-\nabla f_{i}(y)\|^{2}.$
Proof.
Define $F=f_{i}+m\|\cdot\|^{2}/2$ and let $x,y\in\operatorname*{dom}h$ be fixed. Moreover, note that $F$ is convex and satisfies (11) with $L=M+m$. It then follows from Lemma 3.1 with $L=M+m$ that
$\frac{1}{M+m}\|\nabla F(x)-\nabla F(y)\|^{2}$ $=\frac{1}{M+m}\|\nabla f_{i}(x)-\nabla f_{i}(y)+m(x-y)\|^{2}$
$\leq\left\langle\nabla F(x)-\nabla F(y),x-y\right\rangle$
$=\left\langle\nabla f_{i}(x)-\nabla f_{i}(y),x-y\right\rangle+m\|x-y\|^{2}.$
Applying the bound $\|a+b\|^{2}/2\leq\|a\|^{2}+\|b\|^{2}$ with $a=\nabla f_{i}(x)-\nabla f_{i}(y)+m(x-y)$ and b=m(xy)b=-m(x-y), the above inequality then implies
$\left\langle\nabla f_{i}(x)-\nabla f_{i}(y),x-y\right\rangle$ $\geq-m\|x-y\|^{2}+\frac{1}{M+m}\|\nabla f_{i}(x)-\nabla f_{i}(y)+m(x-y)\|^{2}$
$\geq-\left[m+\frac{m^{2}}{2(M+m)}\right]\|x-y\|^{2}+\frac{1}{2(M+m)}\|\nabla f_{i}(x)-\nabla f_{i}(y)\|^{2}.\qed$
The below result gives some technical bounds on changes in the proximal function.
Lemma 3.3.
Given $u,v\in\mathbb{R}^{n}$, let $\psi\in\overline{{\rm Conv}}\,(\mathbb{R}^{n})$ and define
$\Delta:=u-v,\quad\Delta^{p}:=\operatorname{prox}_{\psi}(u)-\operatorname{prox}_{\psi}(v).$
Then, the following statements hold:
  • (a)
    $\|\Delta^{p}\|\leq\left\langle\Delta,\Delta^{p}\right\rangle$;
  • (b)
    $\|\Delta^{p}-\Delta\|^{2}\leq\|\Delta\|^{2}-\|\Delta^{p}\|^{2}$.
Proof.
(a) See [5, Theorem 6.42(a)].
(b) Using part (a), we have that
$\|\Delta^{p}-\Delta\|^{2}$ $=\|\Delta^{p}\|+\|\Delta\|^{2}-2\left\langle\Delta,\Delta^{p}\right\rangle\leq\|\Delta\|^{2}-\|\Delta^{p}\|^{2}.\qed$
We are now ready to present the crucial properties of ${\cal A}_{\lambda}(\cdot)$.
Proposition 3.4.
Define the scalar
$L_{\lambda}=L_{\lambda}(m,M):=\sqrt{1+2\lambda m\left[1+\frac{m}{M+m}\right]}\quad\forall\lambda>0.$ (12)
Then, the following statements hold:
  • (a)
    if $\operatorname*{dom}q_{i}$ is bounded with diameter $C$ for $i\in[k]$, then for any $\lambda>0$ we have
    $\|{\cal A}_{\lambda}(x)-{\cal A}_{\lambda}(y)\|\leq\|x-y\|+2\lambda C\quad\forall x,y\in\operatorname*{dom}h;$ (13)
  • (b)
    if $f_{i}$ satisfies (A2) and $\lambda\leq 1/[2(M+m)]$, then ${\cal A}_{\lambda}(\cdot)$ is $L_{\lambda}$-Lipschitz continuous.
Proof.
(a) Let $x,y\in\operatorname*{dom}h$ and $\lambda>0$ be arbitrary. Moreover, denote $p_{i}(\cdot)=\operatorname{prox}_{q_{i}}(\cdot)$ for $i\in[k]$. Using the definition of ${\cal A}_{\lambda}(\cdot)$, the assumption that $\|p_{i}(z)\|\leq C$ for any $z\in\mathbb{R}^{n}$, and the triangle inequality, we have that
$\|{\cal A}_{\lambda}(x)-{\cal A}_{\lambda}(y)\|$ $=\left\|x-y+\frac{\lambda}{|B|}\sum_{i\in B}[p_{i}(x)-p_{i}(y)]\right\|\leq\|x-y\|+\frac{\lambda}{|B|}\sum_{i\in B}(\|p_{i}(x)\|+\|p_{i}(y)\|)$
$\leq\|x-y\|+2\lambda C.$
(b) Let $x,y,z\in\operatorname*{dom}h$ be arbitrary, and suppose $\lambda\leq 1/(M+m)$. Moreover, denote
$\begin{gathered}\xi(\cdot):={\cal A}_{\lambda}(\cdot,\{f_{i}\},\{q_{i}\}),\quad d:=x-y,\\ \Delta_{i}:=\nabla f_{i}(x)-\nabla f_{i}(y),\quad\Delta_{i}^{p}:=\operatorname{prox}_{q_{i}}(\nabla f_{i}(x))-\operatorname{prox}_{q_{i}}(\nabla f_{i}(y)).\end{gathered}$
Using the previous bound, the fact that $\|\sum_{i=1}^{d}v_{i}\|^{2}\leq d\sum_{i=1}^{d}\|v_{i}\|^{2}$ for any $\{v_{i}\}\subseteq\mathbb{R}^{n}$, and Lemma 3.3(b) with $(\Delta,\Delta^{p})=(\Delta_{i},\Delta_{i}^{p})$, we have that
$\|\xi(x)-\xi(y)\|^{2}$ $=\frac{1}{|B|^{2}}\left\|\sum_{i\in B}\left\{\left[x-\lambda\operatorname{prox}_{q_{i}}(\nabla f_{i}(x))\right]-\left[y-\lambda\operatorname{prox}_{q_{i}}(\nabla f_{i}(y))\right]\right\}\right\|^{2}$
$=\frac{1}{|B|^{2}}\left\|\sum_{i\in B}(d-\lambda\Delta_{i}^{p})\right\|^{2}\leq\frac{1}{|B|}\sum_{i\in B}\left\|d-\lambda\Delta_{i}^{p}\right\|^{2}$
$=\frac{1}{|B|}\sum_{i\in B}\left\|d-\lambda(\Delta_{i}+\Delta_{i}-\Delta_{i}^{p})\right\|^{2}=\frac{2}{|B|}\sum_{i\in B}\left(\|d-\lambda\Delta_{i}\|^{2}+\lambda\|\Delta_{i}-\Delta_{i}^{C}\|^{2}\right)$
$=\frac{2}{|B|}\sum_{i\in B}\left(\|d-\lambda\Delta_{i}\|^{2}+\lambda^{2}\|\Delta_{i}\|^{2}-\lambda^{2}\|\Delta_{i}\|^{2}\right)\leq\frac{2}{|B|}\sum_{i\in B}\left(\|d-\lambda\Delta_{i}\|^{2}+\lambda^{2}\|\Delta_{i}\|^{2}\right)$
$=2\left[\|d\|^{2}+\lambda\left(2\lambda\sum_{i\in B}\left\|\Delta_{i}\right\|^{2}-2\sum_{i\in B}\left\langle d,\Delta_{i}\right\rangle\right)\right].$ (14)
Let us now bound the last two terms of the above expression. Using Lemma 3.1 for the second term, and the fact that $\lambda\leq 1/[2(M+m)]$, we have
$2\lambda\sum_{i\in B}\left\|\Delta_{i}\right\|^{2}-2\sum_{i\in B}\left\langle d,\Delta_{i}\right\rangle\leq 2\lambda\sum_{i\in B}\|\Delta_{i}\|^{2}+2\left[m+\frac{m^{2}}{2(M+m)}\right]\|d\|^{2}-\frac{\sum_{i\in B}\|\Delta_{i}\|^{2}}{(M+m)|B|}$
$=2m\left[1+\frac{m}{2(M+m)}\right]\|d\|^{2}+\left(\underbrace{2\lambda-\frac{1}{M+m}}_{\leq 0}\right)\sum_{i\in B}\|\Delta_{i}\|^{2}\leq 2m\left[1+\frac{m}{2(M+m)}\right]\|d\|^{2}.$ (15)
Using (14), (15), and the definition of $d$, we conclude that
$\|\xi(x)-\xi(y)\|^{2}$ $\leq\|d\|^{2}+\lambda\left(2\lambda\sum_{i\in B}\left\|\Delta_{i}\right\|^{2}-2\sum_{i\in B}\left\langle d,\Delta_{i}\right\rangle\right)$
$\leq\|d\|^{2}+2\lambda m\left[1+\frac{m}{2(M+m)}\right]\|d\|^{2}=L_{\lambda}^{2}\|x-y\|^{2}.\qed$
Before ending this subsection, we give a few remarks about part (b) of the above result. First, $L_{\lambda}(0,M)=1$ and, hence, ${\cal{A}}_{\lambda}(\cdot)$ is nonexpansive when $f_{i}$ is convex for every $i\in[k]$ and $\lambda\leq 1/(2M)$. Second, if $\lambda=1/(2m)$ then $L_{\lambda}(m,0)=\sqrt{3}$ and, hence, ${\cal{A}}_{\lambda}(\cdot)$ $\sqrt{3}$-Lispchitz continuous when $f_{i}$ is concave. Third, like the first remark, $L_{0}(m,M)=1$ implies that ${\cal{A}}_{\lambda}(\cdot)$ is nonexpansive. Finally, when $m=M$ and $\lambda=1/(2m)$, we have $L_{\lambda}(m,M)=\sqrt{5/2}$ and ${\cal{A}}_{\lambda}(\cdot)$ is $\sqrt{5/2}$-Lispchitz continuous.
4 Privacy bounds for DP-SGD
This section develops RDP bounds for the last iterates of DP-SGD under cyclic batch sampling.
For ease of reference, we define
$q_{C}(x)=\begin{cases}1,&\text{if }\|x\|\leq C,\\ \infty,&\text{otherwise},\end{cases}\quad{\rm Clip}_{C}(y)=\operatorname{prox}_{q_{C}}(y).$ (16)
and formally present DP-SGD in Algorithm 1.
Algorithm 1 Cyclically-sampled last-iterate DP-SGD
1:Inputs: $\{f_{i}\}_{i=1}^{k}$, $h$, samples $\{N_{t}\}\subseteq\mathbb{R}^{n}$ from ${\cal N}(0,\sigma^{2}I)$, and $X_{0}\in\operatorname*{dom}h$;
2:Parameters: batch size $b$, stepsize $\lambda$, clipping norm $C$, and iteration limit $T$;
3:Outputs: $X_{T}\in\operatorname*{dom}h$;
4:for $t\leftarrow 1,\ldots,T$ do
5:  $j_{t}\leftarrow b(t-1)\bmod k$
6:  $B_{t}\leftarrow\{j_{t}+1,\ldots,j_{t}+b\}$
7:  $g_{t}\leftarrow(1/b)\sum_{i\in B_{t}}{\rm Clip}_{C}(\nabla f_{i}(X_{t-1}))$
8:  $X_{t}\leftarrow\operatorname{prox}_{\lambda h}\left(X_{t-1}-\lambda g_{t}+N_{t}\right)$
9:return $X_{T}$
Before proceeding, let us give a few remarks about Algorithm 1. First, $j_{t}$ corresponds to the first index of the batch indices $B_{t}$, and if $j_{t}>j_{t-1}$ then $j_{t}-j_{t-1}=b$. Second, $g_{t}$ corresponds to the average of the clipped gradients of $f_{i}$ at $X_{t-1}$. Third, since each $X_{t}$ is generated by a call to $\operatorname{prox}_{\lambda h}(\cdot)$, it holds that $X_{t}\in\operatorname*{dom}h$ for every $t\in[T]$. Finally, notice that the noise $N_{t}$ is added to the displacement $X_{t-1}-\lambda g_{t}$ rather than to $g_{t}$. This setup allows us to develop RDP bounds that scale with both $\lambda$ and $\sigma$.
Now, for the remainder of this section, suppose $h$ satisfies (A1) and let $f_{i}^{\prime}:\mathbb{R}^{n}\mapsto(-\infty,\infty]$ for $i\in[k]$ be such that there exists $i^{*}\in[k]$ where $f_{i}^{\prime}=f_{i}$ for every $i\neq i^{*}$ and $f_{i^{*}}^{\prime}\neq f_{i^{*}}$, i.e., $\{f_{i}\}\sim\{f_{i}^{\prime}\}$. In other words, $i^{*}$ is the index where the neighboring datasets $\{f_{i}\}$ and $\{f_{i}^{\prime}\}$ differ.
For later reference, we also define the assumption:
  • (A3)
    the functions $\{f_{i}^{\prime}\}$ satisfy assumption (A2) with $f_{i}=f_{i}^{\prime}$ for every $i\in[k]$.
The next result shows how the updates in Algorithm 1 are randomized proximal updates applied to the operator $A_{\lambda}(\cdot)$ in (10) with $q_{i}(\cdot)=q_{C}(\cdot)$ from (16).
Lemma 4.1.
Let $\{X_{t}\}$ and $\{X_{t}^{\prime}\}$ denote the iterates generated by Algorithm 1 with functions $\{f_{i}\}$ and $\{f_{i}^{\prime}\}$, respectively, and fixed $\lambda$, $C$, $b$, $\sigma$, $\{N_{t}\}$, $T$, and $X_{0}$ for both sequences of iterates. Moreover, denote
$\phi_{t}(x):={\cal A}_{\lambda}(x,\{f_{i}\},\{q_{C}\}),\quad\phi_{t}^{\prime}(x):={\cal A}_{\lambda}(x,\{f_{i}^{\prime}\},\{q_{C}\}),\quad\forall x\in\operatorname*{dom}h,$
where $q_{C}$ and ${\cal A}_{\lambda}$ are as in (16) and (10), respectively. Then, it holds that
$X_{t}=\operatorname{prox}_{\lambda h}(\phi_{t}(X_{t-1})+N_{t}),\quad X_{t}^{\prime}=\operatorname{prox}_{\lambda h}(\phi_{t}^{\prime}(X_{t-1}^{\prime})+N_{t}),\quad\forall t\geq 1.$
Proof.
This follows immediately from the definition of $\phi_{t}$ and the update rules in Algorithm 1. ∎
We now present some important norm bounds.
Lemma 4.2.
Let $\{X_{t}\}$, $\{X_{t}^{\prime}\}$, $\{\phi_{t}\}$, and $\{\phi_{t}^{\prime}\}$ be as in Lemma 4.1 and denote $\ell=k/b$ and $t^{*}:=\inf_{t\geq 1}\left\{t:i^{*}\in B_{t}\right\}$. Then, it holds that
$\|X_{t^{*}}-X_{t^{*}}^{\prime}\|=0,\quad\|\phi_{s}(x)-\phi_{s}^{\prime}(x)\|\leq\frac{2\lambda C}{b},$ (17)
for every $s\in\{j\ell+t^{*}:j=0,1,$...\} and any $x\in\operatorname*{dom}h$.
Proof.
The identity in (17) follows from the fact that $X_{t}=X_{t}^{\prime}$ for every $t\leq t^{*}$. For the inequality in (17), it suffices to show the bound for $s=t^{*}$ because the batches $B_{t}$ in Algorithm 1 are drawn cyclically. To that end, let $x\in\operatorname*{dom}h$ be fixed. Using the update rule in Algorithm 1, and the fact that $\|{\rm Clip}_{C}(x)\|\leq C$ for every $x\in\mathbb{R}^{n}$, we have that
$\|\phi_{s}(x)-\phi_{s}^{\prime}(x)\|$ $=\frac{1}{b}\left\|\sum_{i\in B_{t^{*}}}\left[x-\lambda{\rm Clip}_{C}(\nabla f_{i^{*}}(x))\right]-\left[x-\lambda{\rm Clip}_{C}(\nabla f_{i^{*}}^{\prime}(x))\right]\right\|$
$=\frac{\lambda}{b}\|{\rm Clip}_{C}(\nabla f_{i^{*}}(x))-{\rm Clip}_{C}(\nabla f_{i^{*}}^{\prime}(x))\|$
$\leq\frac{\lambda}{b}\left[\|{\rm Clip}_{C}(\nabla f_{i^{*}}(x))\|+\|{\rm Clip}_{C}(\nabla f_{i^{*}}^{\prime}(x))\|\right]=\frac{2\lambda C}{b}.\qed$
In the remaining subsections, we show how to use the accumulated results (primarily, Proposition 2.2) to obtain the desired RDP bounds.
4.1 Curvature-independent bounds
This section establishes an RDP bound without assuming any additional conditions on $\{f_{i}\}$. Specifically, using Proposition 2.2(b), we obtain the following bound.
Theorem 4.3.
Let $\{X_{t}\}$, $\{X_{t}^{\prime}\}$, $\lambda$, $b$, $\sigma$, $C$, $\ell$ and $T$ be as in Lemmas 4.1 and 4.2. If $T\geq\ell$, then for any $\lambda>0$ it holds that
$D_{\alpha}(X_{T}\|X_{T}^{\prime})\leq 8\alpha T\left(\frac{\lambda C}{\sigma}\right)^{2}.$ (18)
Proof.
Remark that if $Y_{0}=X_{t^{*}}$, $Y_{0}^{\prime}=X_{t^{*}}^{\prime}$, and $\ell=k/b$, then Proposition 3.4(a), Lemma 4.2, and the fact that $b\geq 1$ imply that the assumptions of Proposition 2.2(b) hold with $\zeta=\bar{s}=2\lambda C$. Consequently, using Proposition 2.2(b) with the previous substitutions, $X_{T}=Y_{T-t^{*}-1}$, $X_{T}^{\prime}=Y_{T-t^{*}-1}^{\prime}$, and $T=T-t^{*}-1$, we have that
$D_{\alpha}(X_{T}\|X_{T}^{\prime})=D_{\alpha}(Y_{T-t^{*}-1}\|Y_{T-t^{*}-1}^{\prime})\leq 8\alpha(T-t^{*}-1)\left(\frac{\lambda C}{\sigma}\right)^{2}\leq 8\alpha T\left(\frac{\lambda C}{\sigma}\right)^{2}.\qed$
We now give two remarks. First, the right-hand-side of the above bound can be made small by decreasing $\lambda$ or increasing $\sigma$. [^6] While decreasing $\lambda$ generally reduces the amount of Gaussian noise added, it can be detrimental to the convergence of the model. This highlights the inherent privacy/accuracy trade-off of DP-SGD [4]. However, the upper bound cannot be made small by increasing the batch size $b$. Second, it improves on the naive bound that is obtained from bounding $X_{T}$ using the triangle inequality and the clipping step. More specifically, using the update rule for $X_{t}$ and the triangle inequality, it is straightforward to show that $\|X_{T}\|\leq\|X_{0}\|+\lambda TC$. Hence, using the Rényi divergence bound between two Gaussian random variables that differ in their mean parameter, one can replace the right-hand-side of (18) with a $\Theta(\alpha[\lambda CT/\sigma]^{2})$ term, which is a factor $T$ worse than the bound in (18).
In the next subsection, we develop improved RDP upper bounds that decrease with increasing $b$, under more restrictive assumptions on $\{f_{i}\}$ and $\{f_{i}^{\prime}\}$.
4.2 Curvature-dependent bounds
This section establishes two RDP bounds under the assumption that $\{f_{i}\}$ and $\{f_{i}^{\prime}\}$ satisfy (A2)–(A3) and $\lambda$ is small relative to $1/(m+M)$.
For the first bound, suppose $\operatorname*{dom}h$ is bounded with diameter
$d_{h}:=\sup\{\|x-y\|:x,y\in\operatorname*{dom}h\}<\infty.$ (19)
Using a special application of Proposition 2.2(a), we obtain the following $d_{h}$-dependent bound.
Theorem 4.4.
Let $\{X_{t}\}$, $\{X_{t}^{\prime}\}$, $\lambda$, $b$, $\sigma$, $C$, and $T$ be as in Lemma 4.1. If $\lambda\leq 1/[2(m+M)]$ and $\{f_{i}\}$ and $\{f_{i}^{\prime}\}$ satisfy (A2)–(A3), then
$D_{\alpha}(X_{T}\|X_{T}^{\prime})$ $\leq\frac{\alpha}{2\sigma^{2}}\left(L_{\lambda}d_{h}+\frac{2\lambda C}{b}\right)^{2},$ (20)
where $L_{\lambda}$ and $d_{h}$ are as in (12) and (19), respectively.
Proof.
Using Proposition 2.2(a) with
$Y_{0}=X_{T-1},\quad Y_{0}^{\prime}=X_{T-1}^{\prime},\quad\tau=d_{h},\quad s=\frac{2\lambda C}{b},\quad L=L_{\lambda},\quad\ell=\frac{k}{b},$
and $E=T=1$, we have that
$D_{\alpha}(X_{T}\|X_{T}^{\prime})$ $\leq D_{\alpha}^{(d_{h})}(X_{T-1}\|X_{T-1}^{\prime})+\frac{\alpha}{2\sigma^{2}}\left(Ld_{h}+\frac{2\lambda C}{b}\right)^{2}\theta_{L}(1)=\frac{\alpha}{2\sigma^{2}}\left(Ld_{h}+\frac{2\lambda C}{b}\right)^{2}.\qed$
Using a special application of Proposition 2.2(b), we present a bound that is linear in terms of the number passes $E$ through the dataset. For conciseness, we leave its proof in Appendix A.3.
Theorem 4.5.
Let $\{X_{t}\}$, $\{X_{t}^{\prime}\}$, $\lambda$, $b$, $C$, and $T$ be as in Lemma 4.1 and denote $E:=\lfloor T/\ell\rfloor$ and $\ell:=k/b$. If $\lambda\leq 1/[2(m+M)]$, $\{f_{i}\}$ and $\{f_{i}^{\prime}\}$ satisfy (A2)–(A3), and $T\geq\ell$, then
$D_{\alpha}(X_{T}\|X_{T}^{\prime})\leq 4\alpha\left(\frac{\lambda C}{b\sigma}\right)^{2}\left[\theta_{L_{\lambda}}(T-E\ell)+E\cdot\theta_{L_{\lambda}}(\ell)\right],$ (21)
where $L_{\lambda}$ is as in (12). Moreover, if $f_{i}$ is convex for every $i\in[k]$, then $m=0$ and
$D_{\alpha}(X_{T}\|X_{T}^{\prime})\leq 4\alpha\left(\frac{\lambda C}{b\sigma}\right)^{2}\left(\max\left\{0,\frac{1}{T-E\ell}\right\}+\frac{E}{\ell}\right).$ (22)
We conclude with a few remarks about the above bounds. First, the bound in (22) is on the same order of magnitude as the bound in [8] in terms of $T$ and $\ell$. However, the right-hand-side of (22) scales linearly with a $\lambda^{2}$ term, which does not appear in [8]. Second, as $\theta_{L_{\lambda}}(\cdot)\leq 1$, the right-hand-side of (21) increases (at most) linearly with respect to the number of dataset passes $E$. Third, substituting $\sigma=\Theta({C/[b\sqrt{\epsilon}]})$ in (21) yields a bound that depends linearly on $\varepsilon$ and i invariant to changes in $C$ and $b$.
In Appendix C, we discuss further choices of the parameters in (21) and their implications. Some highlights are as follows. First, when $\ell$ is sufficiently large, the terms in (21) that involve $\theta_{L_{\lambda}}(\ell)$ become $\Theta(1+Em/[M+m])$, which shows that the effect of $E$ diminishes as $m\to 0$ (with the caveat that $\ell$ grows as well). Second, when $\lambda\leq\min\{1/\sqrt{E},1/2[m+M]\}$, the resulting bound is independent of $E$.
Appendix A Technical proofs
A.1 Proof of Lemma 2.1
The below result gives two calculus rules for the shifted Rényi divergence given in (3). In particular, the proof of the second rule is a minor modification of the proof given for [8, Lemma 21].
Lemma A.1.
For random variables $\{X^{\prime},X,Z\}$ and $a,s\geq 0$ and $\alpha\in(1,\infty)$, it holds that
  • (a)
    $D_{\alpha}^{(\tau)}(X+Z\|X^{\prime}+Z)\leq D_{\alpha}^{(\tau+a)}(X\|X^{\prime})+\sup_{c\in\mathbb{R}^{n}}\{D_{\alpha}([Z+c]\|Z):\|c\|\leq a\}$;
  • (b)
    for some $L,\zeta>0$, if $\phi^{\prime}$ and $\phi$ satisfy
    $\sup_{u}\|\phi^{\prime}(u)-\phi(u)\|\leq s,\quad\|\Phi(x)-\Phi(y)\|\leq L\|x-y\|+\zeta,\quad\forall\Phi\in\{\phi,\phi^{\prime}\},$
    for any $x,y\in\operatorname*{dom}\phi\cap\operatorname*{dom}\phi^{\prime}$, then
    $D_{\alpha}^{(L\tau+\zeta+s)}(\phi(X)\|\phi^{\prime}(X^{\prime}))\leq D_{\alpha}^{(\tau)}(X\|X^{\prime}).$
Proof.
(a) See [8, Lemma 20].
(b) By the definitions of of $D_{\alpha}^{(\tau)}(\mu\|\nu)$ and ${\cal W}_{\infty}(\mu,\nu)$, there exist a joint distribution $(X,Y)$ such that $D_{\alpha}(Y\|X^{\prime})=D_{\alpha}^{(\tau)}(X\|X^{\prime})$ and $\|X-Y\|\leq\tau$ almost surely. Now, the post-processing property of Rényi divergence implies that
$D_{\alpha}(\phi(Y)\|\phi^{\prime}(X^{\prime}))\leq D_{\alpha}(\phi(Y)\|X^{\prime})\leq D_{\alpha}(Y\|X^{\prime})=D_{\alpha}^{(\tau)}(X\|X^{\prime}).$
Using our assumptions on $\phi$ and $\phi^{\prime}$ and the triangle inequality, we then have
$\|\phi(X)-\phi^{\prime}(Y)\|$ $\leq\|\phi(X)-\phi(Y)\|+\|\phi(Y)-\phi^{\prime}(Y)\|$
$\leq L\|X-Y\|+\zeta+s\leq L\tau+\zeta+s,$
almost surely. Combining the previous two inequalities, yields the desired bound in view of the definitions of $D_{\alpha}^{(\tau)}(\mu\|\nu)$ and ${\cal W}_{\infty}(\mu,\nu)$. ∎
We are now ready to give the main proof of this subsection.
Proof of Lemma 2.1..
We first have that
$\sup_{\tau\in\mathbb{R}^{n}}\{D_{\alpha}([Z+c]\|Z):\|c\|\leq a\}=\sup_{c\in\mathbb{R}^{n}}\left\{\frac{\alpha c^{2}}{2\sigma^{2}}:\|c\|\leq a\right\}=\frac{\alpha a^{2}}{2\sigma^{2}},$ (23)
from the well-known properties of the Rényi divergence. Using (23), Lemma A.1(a) with $(X,X^{\prime})=(\phi(Y),\phi^{\prime}(Y^{\prime}))$, and Lemma A.1(b) with $(\phi,\phi^{\prime},L,s)\in\{(\phi,\phi^{\prime},L,s),({\rm prox}_{\psi},{\rm prox}_{\psi},1,0)\}$, we have
$D_{\alpha}^{(L\tau+s-a)}({\rm prox}_{\psi}(\phi(Y)+Z)\|{\rm prox}_{\psi}(\phi^{\prime}(Y^{\prime})+Z))\leq D_{\alpha}^{(L\tau+s-a)}(\phi(Y)+Z\|\phi^{\prime}(Y^{\prime})+Z)$
$\leq D_{\alpha}^{(L\tau+s)}(\phi(Y)\|\phi^{\prime}(Y^{\prime}))+\frac{\alpha a^{2}}{2\sigma^{2}}\leq D_{\alpha}^{(\tau)}(Y\|Y^{\prime})+\frac{\alpha a^{2}}{2\sigma^{2}}.\qed$
A.2 Proof of Proposition 2.2
We start with a technical lemma.
Lemma A.2.
Given scalars $L>1$ and positive integer $T\geq 1$, let
$S_{T}:=\sum_{i=0}^{T-1}L^{2i},\quad b_{t}:=\left(\frac{L^{T-t}}{S_{T}}\right)L^{T-1},\quad R_{t}:=L^{t-1}-\sum_{i=1}^{t}b_{i}L^{t-i},\quad t\geq 0$ (24)
Then, for every $t\in[T]$,
  • (a)
    $R_{t+1}=LR_{t}-b_{t+1}$;
  • (b)
    $R_{t}\geq 0$ and $R_{T}=0$;
  • (c)
    $\sum_{t=1}^{T}b_{t}^{2}=\theta_{L}(T)$.
Proof.
Let $t\in[T]$ be fixed.
(a) This is immediate from the definition of $R_{t}$.
(b) We have that
$S_{T}R_{t}$ $=S_{T}\left(L^{t}-\sum_{i=1}^{t}b_{i}L^{t-i}\right)=\sum_{i=0}^{T-1}L^{2i+t-1}-\sum_{i=1}^{t}L^{2T+t-2i-1}=L^{t-1}\left[\sum_{i=0}^{T-1}L^{2i}-\sum_{i=1}^{t}L^{2(T-i)}\right]$
$=L^{t-1}\sum_{i=0}^{T-1-t}L^{2i}\geq 0.$
Evaluating the above expression at $t=T$ clearly gives $R_{T}=0$.
(c) The case of $T=0$ is immediate. For the case of $T\geq 1$, we use the definitions of $b_{t}$ and $S_{T}$ to obtain
$\sum_{t=1}^{T}b_{t}^{2}$ $=\frac{L^{2(T-1)}\sum_{i=0}^{T-1}L^{2i}}{\left(\sum_{i=0}^{T-1}L^{2i}\right)^{2}}=\frac{L^{2(T-1)}}{S_{T}}=\theta_{L}(T).\qed$
The next result considers the case of a single dataset pass with $T=\ell$.
Lemma A.3.
Let $L$, $\tau$, $\zeta$, $\bar{s}$, $\{Y_{t}\}$, $\{Y_{t}^{\prime}\}$, $\ell$, and $T$ be as in Proposition 2.2. If $T=\ell$, then the following statements hold:
  • (a)
    if $\zeta=0$, then
    $D_{\alpha}(Y_{T}\|Y_{T}^{\prime})-D_{\alpha}^{(\tau)}(Y_{0}\|Y_{0}^{\prime})\leq\frac{\alpha}{2}\left(\frac{L\tau+\bar{s}}{\sigma}\right)^{2}\theta_{L}(T);$ (25)
  • (b)
    if $\tau=0$, $L=1$, and $\zeta=\bar{s}$, then
    $D_{\alpha}(Y_{T}\|Y_{T}^{\prime})-D_{\alpha}(Y_{0}\|Y_{0}^{\prime})\leq 2\alpha T\left(\frac{\zeta}{\sigma}\right)^{2}$ (26)
Proof.
(a) Let $s=\bar{s}$. Our goal is to recursively apply (5) with suitable choices of the free parameter $a$ at each application. Specifically, let $\{(b_{t},R_{t},S_{T})\}$ be as in (24), and define
$a_{t}:=(L\tau+s)b_{t}\quad\forall t\geq 1.$
Using Lemma A.2(a)–(b), we first have $L\tau+s-a_{1}=(L\tau+s)R_{1}\geq 0$ and, hence, by Lemma 2.1, we have
$D_{\alpha}^{([L\tau+s]R_{1})}(Y_{1}\|Y_{1}^{\prime})=D_{\alpha}^{(L\tau+{s}-a_{1})}(Y_{1}\|Y_{1}^{\prime})\leq D_{\alpha}^{(\tau)}(Y_{0}\|Y_{0}^{\prime})+\frac{\alpha a_{1}^{2}}{2\sigma^{2}}.$
Since Lemma A.2(a)–(b) also implies $R_{t}\geq 0$ and we have $s_{t}=0$ for $t\geq 2$, we repeatedly apply Lemma 2.1 with $(a,\tau)=(a_{t},\tau_{t})=(a_{t},0)$ for $t\geq 2$ to obtain
$D_{\alpha}^{(\tau)}(Y_{0}\|Y_{0}^{\prime})$ $\geq D_{\alpha}^{([L\tau+s]R_{1})}(Y_{1}\|Y_{1}^{\prime})-\frac{\alpha a_{1}^{2}}{2\sigma^{2}}\geq D_{\alpha}^{([L\tau+s]LR_{1}-a_{2})}(Y_{2}\|Y_{2}^{\prime})-\frac{\alpha(a_{1}^{2}+a_{2}^{2})}{2\sigma^{2}}$
$=D_{\alpha}^{([L\tau+s]R_{2})}(Y_{2}\|Y_{2}^{\prime})-\frac{\alpha(a_{1}^{2}+a_{2}^{2})}{2\sigma^{2}}\geq\cdots$
$\geq D_{\alpha}^{([L\tau+s]R_{T})}(Y_{T}\|Y_{T}^{\prime})-\frac{\alpha\sum_{i=1}^{T}a_{i}^{2}}{2\sigma^{2}}=D_{\alpha}^{(0)}(Y_{T}\|Y_{T}^{\prime})-\frac{\alpha\sum_{i=1}^{T}a_{i}^{2}}{2\sigma^{2}}$
$=D_{\alpha}(Y_{T}\|Y_{T}^{\prime})-\frac{\alpha\sum_{i=1}^{T}a_{i}^{2}}{2\sigma^{2}}.$
It now remains to bound $\alpha\sum_{i=1}^{T}a_{i}^{2}/(2\sigma^{2})$. Using Lemma A.2(c) and the fact that $T=\ell$ and $\bar{s}=s$, we have
$\frac{\alpha\sum_{i=1}^{T}a_{i}^{2}}{2\sigma^{2}}$ $=\frac{\alpha}{2\sigma^{2}}\left[(L\tau+s)^{2}\sum_{i=2}^{T}b_{i}^{2}\right]\leq\frac{\alpha}{2}\left(\frac{L\tau+\bar{s}}{\sigma}\right)^{2}\theta_{L}(T).$
Combining this bound with the previous one yields the desired conclusion.
(b) Let $s=\bar{s}$. Similar to (a), our goal is to recursively apply (5) with suitable choices of the free parameter $a$ at each application. For this setting, let $a_{1}=\zeta+s$ and $a_{t}=\zeta$ for $t\geq 2$. Using the fact that $\tau=0$ and $L=1$ and Lemma 2.1, we first have that
$D_{\alpha}(Y_{1}\|Y_{1}^{\prime})=D_{\alpha}^{(0)}(Y_{1}\|Y_{1}^{\prime})=D_{\alpha}^{(s+\zeta-a_{1})}(Y_{1}\|Y_{1}^{\prime})\leq D_{\alpha}^{(0)}(Y_{0}\|Y_{0}^{\prime})+\frac{\alpha a_{1}^{2}}{2\sigma^{2}}.$
We then repeatedly apply Lemma 2.1 with $(a,\tau)=(a_{t},0)$ for $t\geq 2$ to obtain
$D_{\alpha}^{(0)}(Y_{0}\|Y_{0}^{\prime})$ $\geq D_{\alpha}^{(0)}(Y_{1}\|Y_{1}^{\prime})-\frac{\alpha a_{1}^{2}}{2\sigma^{2}}\geq D_{\alpha}^{(\zeta-a_{2})}(Y_{2}\|Y_{2}^{\prime})-\frac{\alpha(a_{1}^{2}+a_{2}^{2})}{2\sigma^{2}}$
$=D_{\alpha}^{(0)}(Y_{2}\|Y_{2}^{\prime})-\frac{\alpha(a_{1}^{2}+a_{2}^{2})}{2\sigma^{2}}\geq\cdots$
$\geq D_{\alpha}^{(\zeta-a_{T})}(Y_{T}\|Y_{T}^{\prime})-\frac{\alpha\sum_{i=1}^{T}a_{i}^{2}}{2\sigma^{2}}=D_{\alpha}^{(0)}(Y_{T}\|Y_{T}^{\prime})-\frac{\alpha\sum_{i=1}^{T}a_{i}^{2}}{2\sigma^{2}}$
$=D_{\alpha}(Y_{T}\|Y_{T}^{\prime})-\frac{\alpha\sum_{i=1}^{T}a_{i}^{2}}{2\sigma^{2}}.$
It now remains to bound $\alpha\sum_{i=1}^{T}a_{i}^{2}/(2\sigma^{2})$. Using the definition of $\{a_{t}\}$ and the fact that $\zeta=s$, it holds that
$\frac{\alpha\sum_{i=1}^{T}a_{i}^{2}}{2\sigma^{2}}\leq\frac{\alpha}{2\sigma^{2}}\left[4\zeta^{2}+(T-1)\zeta^{2}\right]\leq 2\alpha T\left(\frac{\zeta}{\sigma}\right)^{2}.$
Combining this bound with the previous one yields the desired conclusion. ∎
We are now ready to give the main proof of this subsection.
Proof of Proposition 2.2.
(a) Let $s=\bar{s}$. For convenience, define
${\cal B}_{1}(\tau,T)$ $:=\frac{\alpha}{2}\left(\frac{L\tau+s}{\sigma}\right)^{2}\theta_{L}(T),$
${\cal B}_{2}$ $:=\frac{\alpha}{\sigma^{2}}\left[(L\tau+s)^{2}+s^{2}\left\{(E-1)\theta_{L}(\ell)+\theta_{L}(T-E\ell)\right\}\right].$
Using Lemma A.3(a), we have that for the first $\ell$ iterates,
$D_{\alpha}(Y_{\ell}\|Y_{\ell}^{\prime})-D_{\alpha}^{(\tau)}(Y_{0}\|Y_{0}^{\prime})\leq{\cal B}_{1}(\tau,\ell).$
Similarly, using part Lemma A.3(a) with $\tau=0$, we have that
$D_{\alpha}(Y_{[k+1]\ell}\|Y_{[k+1]\ell}^{\prime})-D_{\alpha}^{(0)}(Y_{\ell}\|Y_{\ell}^{\prime})\leq{\cal B}_{1}(0,\ell),$
for any $1\leq k\leq E-1$. Finally, using part Lemma A.3(a) with $T=T-E\ell$ and $\tau=0$, we have that
$D_{\alpha}(Y_{T}\|Y_{T}^{\prime})-D_{\alpha}^{(0)}(Y_{E\ell}\|Y_{E\ell}^{\prime})\leq{\cal B}_{1}(0,T-E\ell).$
Summing the above three inequalities, using the fact that $D_{\alpha}^{(0)}(X\|Y)=D_{\alpha}(X\|Y)$, and using the definition of ${\cal B}_{2}$ we conclude that
$D_{\alpha}(Y_{T}\|Y_{T}^{\prime})-D_{\alpha}(Y_{0}\|Y_{0}^{\prime})$
$\leq{\cal B}_{1}(\tau,\ell)+(E-1){\cal B}_{1}(0,\ell)+{\cal B}_{1}(0,T-E\ell)={\cal B}_{2}.$
(b) The proof follows similarly to (a). Repeatedly using Lemma A.3(b) at increments of $\ell$ iterations up to iteration $E\ell$, we have that
$D_{\alpha}(Y_{E\ell}\|Y_{E\ell}^{\prime})$ $\leq D_{\alpha}(Y_{(E-1)\ell}\|Y_{(E-1)\ell}^{\prime})+2\alpha\ell\left(\frac{\zeta}{\sigma}\right)^{2}\leq D_{\alpha}(Y_{(E-2)\ell}\|Y_{(E-2)\ell}^{\prime})+4\alpha\ell\left(\frac{\zeta}{\sigma}\right)^{2}\leq\cdots$
$\leq D_{\alpha}(Y_{0}\|Y_{0}^{\prime})+2\alpha E\ell\left(\frac{\zeta}{\sigma}\right)^{2}.$
For the last $T-E\ell$ iterations, we use Lemma A.3(b) with $T=T-E\ell$ and the above bound to obtain
$D_{\alpha}(Y_{T}\|Y_{T}^{\prime})$ $\leq D_{\alpha}(Y_{E\ell}\|Y_{E\ell}^{\prime})+2\alpha[T-E\ell]\left(\frac{\zeta}{\sigma}\right)^{2}\leq D_{\alpha}(Y_{0}\|Y_{0}^{\prime})+2\alpha T\left(\frac{\zeta}{\sigma}\right)^{2}.\qed$
A.3 Proof of Theorem 4.5
The below result gives a general divergence bound that primarily depends on $i^{*}$, $E$, and $\ell$.
Proposition A.4.
Let $\{X_{t}\}$, $\{X_{t}^{\prime}\}$, $\lambda$, $b$, $C$, and $T$ be as in Theorem 4.4, and let $t^{*}:=\inf_{t\geq 1}\left\{t:i^{*}\in B_{t}\right\}$. If $\lambda\leq 1/[2(m+M)]$ and $T\geq t^{*}+1$, then
$D_{\alpha}(X_{T}\|X_{T}^{\prime})$ $\leq 2\alpha\left(\frac{\lambda C}{b\sigma}\right)^{2}\left[E\theta_{L_{\lambda}}(\ell)+\theta_{L_{\lambda}}(T-t^{*}-1-E\ell)\right],$ (27)
where $E:=\lfloor(T-t^{*}-1)/\ell\rfloor$, $\ell:=k/b$, and $L_{\lambda}$ is as in (12).
Proof.
Using Lemma 4.2(b) and Proposition 2.2(a) with
$Y_{0}=X_{t^{*}},\quad Y_{0}^{\prime}=X_{t^{*}}^{\prime},\quad\tau=0,\quad s=\frac{2\lambda C}{b},\quad L=L_{\lambda},\quad\ell=\frac{k}{b},$
and $T=T-t^{*}-1$, we have that
$D_{\alpha}(X_{T}\|X_{T}^{\prime})$ $\leq D_{\alpha}^{(0)}(X_{0}\|X_{0})+2\alpha\left(\frac{\lambda C}{b\sigma}\right)^{2}\left[E\theta_{L_{\lambda}}(\ell)+\theta_{L_{\lambda}}(T-t^{*}-1-E\ell)\right]$
$=2\alpha\left(\frac{\lambda C}{b\sigma}\right)^{2}\left[E\theta_{L_{\lambda}}(\ell)+\theta_{L_{\lambda}}(T-t^{*}-1-E\ell)\right].\qed$
We are now ready to give the main proof of this subsection.
Proof of Theorem 4.5.
The bound in (21) follows from Proposition A.4, the fact that $t^{*}\geq 1$, and the fact that $\theta_{L_{\lambda}}(s)\leq 1$ for every $s\geq 1$, and the fact that $\theta_{L_{\lambda}}(s)$ is monontonically decreasing for $s\geq 1$. Now, assume that each $f_{i}$ is convex. The fact that $m=0$ is well-known. Next, observe that the definition of $L_{\lambda}=L_{\lambda}(m,M)$ in (12) implies that $L_{\lambda}(0,M)=1$ and the definition of $\theta_{L}(\cdot)$ in (6) implies $\theta_{1}(s)=1/s$ for every $s\geq 1$ and $\theta_{1}(0)=0$. The bound in (22) now follows from using the previous identities for $\theta_{1}(\cdot)$ in (21). ∎
Appendix B Choice of residuals
This appendix briefly discusses the choice of residuals $\{a_{t}\}$ that are used in the proof of Proposition 2.2(a) through Lemma 2.1.
In the setup of Proposition 2.2(a), it is straightforward to show that if $\{a_{t}\}$ is a nonnegative sequence of scalars such that
$\tilde{R}_{t}:=L^{t-1}(L\tau+s)-\sum_{i=1}^{t}a_{i}L^{t-i}\geq 0,\quad\tilde{R}_{T}=0,$
then repeatedly applying Lemma 2.1 with $a=a_{t}$ yields
$D_{\alpha}(Y_{T}\|Y_{T}^{\prime})-D_{\alpha}^{(\tau)}(Y_{0}\|Y_{0}^{\prime})\leq\frac{\alpha}{2\sigma^{2}}\sum_{i=1}^{T}a_{i}^{2}.$ (28)
Hence, to obtain the tightest bound of the form in (28), we need to solve the quadratic program
$(P)\quad\min\$$\frac{1}{2}\sum_{i=1}^{T}a_{i}^{2}$
s.t $\tilde{R}_{t}\geq 0\quad\forall t\in[T-1],$
$\tilde{R}_{T}=0.$
If we ignore the inequality constraints, the first order optimality condition of the resulting problem is that there exists $\xi\in\mathbb{R}$ such that
$a_{i}=\xi L^{t-i}\quad\forall t\in[T],\quad\tilde{R}_{T}=0.$
The latter identity implies that
$L^{T-1}(L\tau+s)=\xi\sum_{i=1}^{T}L^{2(T-i)}=\xi\sum_{i=0}^{T-1}L^{2i}$
which then implies
$a_{i}=\frac{L^{T-1}(L\tau+s)L^{t-i}}{\sum_{i=0}^{T-1}L^{2i}}\quad\forall t\in[T].$
Hence, to verify that the above choice of $a_{i}$ is optimal for $(P)$, it remains to verify that $\tilde{R}_{t}\geq 0$ for $t\in[T-1]$. Indeed, this follows from Lemma A.2(b) after normalizing for the $L\tau+s$ factor. As a consequence, the right-hand-side of (28) is minimized for our choice of $a_{i}$ above.
Appendix C Parameter choices
Let us now consider some interesting values for $\lambda$, $\sigma$, and $\ell$.
The result below establishes a useful bound on $\theta_{L}(s)$ for sufficiently large enough values of $s$.
Lemma C.1.
For any $L>1$ and $\xi>1$, if $s\geq\log_{L}\sqrt{\xi/(\xi-1)}$ then $\theta_{L}(s)\leq\xi\left(1-L^{-2}\right).$
Proof.
Using the definition of $\theta_{L}(\cdot)$, we have
$\theta_{L}(s)$ $=\frac{L^{-2(s-1)}}{\sum_{i=0}^{s-1}L^{2i}}=\frac{L^{2s}-L^{2(s-1)}}{L^{2s}-1}=\frac{1-L^{-2}}{1-L^{-2s}}\leq\frac{1-L^{-2}}{1-L^{-2\log_{L}\sqrt{\xi/(\xi-1)}}}$
$=\frac{1-L^{-2}}{1-(\xi-1)/\xi}=\xi\left(1-\frac{1}{L^{2}}\right).\qed$
Corollary C.2.
Let $\alpha>1$ and $\varepsilon>0$ be fixed, and let $\{X_{t}\}$, $\{X_{t}^{\prime}\}$, $b$, $C$, $E$, $\ell$, $\lambda$, and $T$ be as in Theorem 4.5. Moreover, define
$\overline{\lambda}(\rho):=\frac{1}{2(M+\rho)},\quad\overline{\sigma}_{\varepsilon}(\rho):=\frac{C\cdot\overline{\lambda}(\rho)}{2b}\sqrt{\frac{1}{\alpha\varepsilon}\left(1+\left[\frac{4\rho}{M+\rho}\right]E\right)},\quad\overline{\ell}(\rho):=\frac{\log 2}{\log\left[1+\rho\overline{\lambda}(\rho)\right]},$
for every $\rho>0$. If $\lambda=\overline{\lambda}(m)$, $\sigma\geq\overline{\sigma}_{\varepsilon}(m)$, and $\ell\geq\overline{\ell}(m)$, then
$D_{\alpha}(X_{T}\|X_{T}^{\prime})\leq 4\alpha\left[\frac{C\cdot\overline{\lambda}(m)}{b\cdot\overline{\sigma}_{\varepsilon}(m)}\right]^{2}\left[1+\frac{4m}{M+m}\right],$
and the corresponding instance of Algorithm 1 is $(\alpha,\varepsilon)$-Rényi-DP.
Proof.
For ease of notation, denote $\overline{\lambda}=\overline{\lambda}(m)$, $L=L_{\overline{\lambda}}$, $\overline{\sigma}=\overline{\sigma}(m)$, and $\overline{\ell}=\overline{\ell}(m)$. We first note that
$L=L_{\overline{\lambda}}=\sqrt{1+\frac{m}{M+m}\left[1+\frac{m}{M+m}\right]}\geq\sqrt{1+m\overline{\lambda}(m)},$
which implies
$\ell\geq\overline{\ell}=\frac{\log\sqrt{2}}{\log\sqrt{1+m\overline{\lambda}}}=\frac{\log\sqrt{2}}{\log L}=\log_{L}\sqrt{2}.$
Consequently, using Lemma C.1 with $(\xi,s)=(2,\ell)$ and the definitions of $L_{\lambda}(\cdot)$ and $\overline{\lambda}(\cdot)$, we have that
$\theta_{L}(\ell)\leq 2\left(1-\frac{1}{L^{2}}\right)=2\left(\frac{2m}{2(M+m)}\left[1+\frac{m}{M+m}\right]\right)\leq\frac{4m}{M+m}.$
Using the above bound and Theorem 4.5 with $(\lambda,\sigma,L)=(\overline{\lambda},\overline{\sigma},L_{\overline{\lambda}})$, we obtain
$D_{\alpha}(X_{T}\|X_{T}^{\prime})$ $\leq 4\alpha\left(\frac{\overline{\lambda}C}{b\sigma}\right)^{2}\left[1+E\theta_{L}(\ell)\right]\leq 4\alpha\left(\frac{\overline{\lambda}C}{b\sigma}\right)^{2}\left[1+\frac{4Em}{M+m}\right]$
$\leq 4\alpha\left(\frac{\overline{\lambda}C}{b\overline{\sigma}}\right)^{2}\left[1+\frac{4Em}{M+m}\right]\leq\varepsilon.$
In view of the fact that Algorithm 1 returns the last iterate $X_{T}$ (or $X_{T}^{\prime}$), the conclusion follows. ∎
Some remarks about Corollary C.2 are in order. First, $\sigma_{\varepsilon}^{2}(m)$ increases linearly with the number of dataset passes $E$. Second, the smaller $m$ is the smaller the effect of $E$ on $\sigma_{\varepsilon}(m)$ is. Fourth, $\lim_{m\to 0}\overline{\ell}(m)=\infty$ which implies that the reducing the dependence on $E$ in $\sigma_{\varepsilon}(m)$ leads to more restrictive choices on $\ell$. Finally, it is worth emphasizing that the restrictions on $\ell$ can be removed by using (21) directly. However, the resulting bounds are less informative in terms of the topological constants $m$ and $M$.
We now present an RDP bound that is independent of $E$ when $\lambda$ is sufficiently small.
Corollary C.3.
Let $\{X_{t}\}$, $\{X_{t}^{\prime}\}$, $b$, $C$, $E$, $\ell$, $\lambda$, and $T$ be as in Theorem 4.5. If
$\lambda\leq\min\left\{\frac{1}{\sqrt{E}},\frac{1}{2(m+M)}\right\},$
then we have
$D_{\alpha}(X_{T}\|X_{T}^{\prime})\leq 4\alpha\left(\frac{C}{b\sigma}\right)^{2}\left[1+\theta_{L_{\lambda}}(\ell)\right].$
Proof.
Using Theorem 4.5 and the fact that $\theta_{L}(\cdot)\leq 1$ for any $L>1$, we have that
$D_{\alpha}(X_{T}\|X_{T}^{\prime})$ $\leq 4\alpha\left(\frac{\lambda C}{b\sigma}\right)^{2}\left[\theta_{L_{\lambda}}(T-E\ell)+E\theta_{L_{\lambda}}(\ell)\right]\leq 4\alpha\left(\frac{\lambda C}{b\sigma}\right)^{2}\left[1+E\theta_{L_{\lambda}}(\ell)\right]$
$\leq 4\alpha\left(\frac{C}{b\sigma}\right)^{2}\left[\frac{1}{E}+\theta_{L_{\lambda}}(\ell)\right]\leq 4\alpha\left(\frac{C}{b\sigma}\right)^{2}\left[1+\theta_{L_{\lambda}}(\ell)\right].\qed$